# DBLP: Noise Bridge Consistency Distillation For Efficient And Reliable Adversarial Purification

### v1

@@ Introduction
where the thresholds are calculated using Otsu (Otsu 1979) algorithm.
@@
Experiments
Experimental Settings
Datasets
We conduct extensive experiments to validate the effectiveness and efficiency of our proposed method across several widely-used datasets, including CIFAR-10 (Krizhevsky et al. 2009), ImageNet (Deng et al. 2009), and CelebA (Liu et al. 2015). CIFAR-10 consists of 60,000 color images of size 32×32 across 10 object classes, representing general-purpose natural scenes. ImageNet is a large-scale visual database with over 14 million human-annotated images spanning more than 20,000 categories. CelebA contains over 200,000 celebrity face images, each annotated with 40 facial attributes and five landmark points.
Training Settings
For our pretrained diffusion backbone, we use Stable Diffusion v1.5 (Rombach et al. 2022). The distillation process is trained for 20,000 iterations with a batch size of 4, a learning rate of 8e-6, and a 500-step warm-up schedule. For our leapfrog solver $\Psi_{\mathrm{leap}}$, we set $k=20$ in Equation (13) and $h=0.8$ in Equation (15). During training, adversarial noise is generated using PGD-100 with $\epsilon=4/255$, targeting a ResNet-50 (He et al. 2016) classifier.
Evaluation Metrics
We evaluate our approach using multiple metrics: clean accuracy (performance on clean data), robust accuracy (performance under adversarial attack), inference time, and image quality metrics including LPIPS (Zhang et al. 2018), PSNR, and SSIM (Horé and Ziou 2010).
@@
Method Type Attack Standard Acc. Robust Acc. Architecture
w/o Defense -- PGD-100 80.55 0.01 Res-50
(Schott et al. 2019) AT PGD-40 72.70 47.00 Res-152
(Wang et al. 2020) PGD-100 53.83 28.04 Res-50
ConvStem (Singh et al. 2023) AutoAttack 77.00 57.70 ConvNeXt-L
MeanSparse (Amini et al. 2024) AutoAttack 77.96 59.64 ConvNeXt-L
DiffPure (Nie et al. 2022) AP PGD-100 68.22 42.88 Res-50
DiffPure (Nie et al. 2022) AutoAttack 71.16 44.39 WRN-50-2
(Bai et al. 2024) PGD-200 ($\epsilon=8/255$) 70.41 41.70 Res-50
(Lee and Kim 2023) PGD+EOT 70.74 42.15 Res-50
(Lin et al. 2025) PGD+EOT 68.75 45.90 Res-50
(Zollicoffer et al. 2025) PGD-200 ($\epsilon=8/255$) 73.98 56.54 Res-50
MimicDiffusion (Song et al. 2024) AutoAttack 66.92 61.53 Res-50
ScoreOpt (Zhang et al. 2023a) Transfer-PGD 71.68 62.10 WRN-50-2
(Pei et al. 2025) PGD-200 ($\epsilon=8/255$) 77.15 65.04 Res-50
OSCP (Lei et al. 2025) PGD-100 77.63 73.89 Res-50
DBLP (Ours) PGD-100 78.2 75.6 Res-50
DBLP (Ours) AutoAttack 78.0 74.8 Res-50
DBLP (Ours) PGD-200 ($\epsilon=8/255$) 77.4 74.2 Res-50
[…]
Results
CIFAR-10
We first conduct experiments on the CIFAR-10 dataset, evaluating our method under adversarial threats constrained by $\ell_{\infty}$, $\ell_{1}$, and $\ell_{2}$ norms. Since DBLP is trained under $\ell_{\infty}$ attacks, this scenario is considered a seen threat, while the $\ell_{1}$ and $\ell_{2}$ settings are treated as unseen threats. The results are presented in Table 1. Although DBLP belongs to the category of adversarial purification methods, its access to the victim classifier makes it comparable to SOTA adversarial training and purification methods. Adversarial training performs well on seen threats but generalizes poorly to unseen ones, and DiffPure variants offer limited gains. In contrast, DBLP achieves substantially higher robust accuracy on both seen and unseen threats, while preserving strong clean accuracy. It outperforms prior methods by XXXX%, highlighting its robustness, generalization, and efficiency.
@@ Results
ImageNet
Method runtime (s)
GDMP (Wang et al. 2022) 19$\sim$ 19∼ 19
DiffPure (Nie et al. 2022) 25$\sim$ 25∼ 25
OSCP (Lei et al. 2025) 0.8$\sim$ 0.8∼ 0.8
[…]
Celeb-A
We further validated the effectiveness of our method on a subset of the CelebA-HQ dataset by evaluating it against three representative victim models: ArcFace (AF) (Deng et al. 2019), FaceNet (FN) (Schroff, Kalenichenko, and Philbin 2015), and MobileFaceNet (MFN) (Chen et al. 2018). Leveraging model weights pretrained on ImageNet, we applied our purification framework to adversarial face images. As shown in Table 2, DBLP significantly enhances purification performance on facial data, demonstrating its robust generalization across image resolutions and domains.
Transferability
We further evaluated DBLP under the Diff-PGD attack (Xue et al. 2023). The LCM was trained using PGD-generated adversarial noise on ResNet-50 and tested for transfer robustness across diverse architectures, including ResNet-50/152, WideResNet-50-2 (Zagoruyko and Komodakis 2016), ConvNeXt-B (Liu et al. 2022), ViT-B-16 (Kolesnikov et al. 2021), and Swin-B (Liu et al. 2021). As shown in Table 6, DBLP consistently outperforms prior SOTA methods under Diff-PGD-10, demonstrating strong cross-architecture robustness.
DiffPure (Nie et al. 2022) 53.8 49.4 52.2 42.9 16.6 45.1
OSCP (Lei et al. 2025) 59.0 56.5 57.9 49.1 34.1 53.9
DBLP (Ours) 63.0 59.4 60.7 52.4 38.2 58.3
[…]
Robust Acc. LPIPS $\downarrow$ SSIM $\uparrow$
w/o Edge Map 74.2 0.1386 0.7409
Edge Map 74.8 0.1172 0.7430
Pyramid Edge Map 75.6 0.1012 0.7655
Table 7: Ablation study of adaptive semantic enhancement.
Image Quality
Beyond correct classification, adversarial purification also seeks to maintain visual fidelity relative to the clean input. As shown in Table 4, DBLP achieves strong image quality across all three metrics, with purified outputs $\mathbf{x}{\mathrm{pur}}$ closely matching both adversarial $\mathbf{x}{\mathrm{adv}}$ and clean images $\mathbf{x}$. This highlights DBLP’s superior visual quality. Qualitative results in Figure 2 further confirm its ability to preserve fine-grained details.
Inference Time
@@
Ablation Study
We conduct ablation studies to assess the adaptive semantic enhancement module in DBLP, with results in Table 7. Omitting edge maps leads to a small drop in robust accuracy but a significant decline in image quality. Using pyramid edge maps further improves both metrics, showing that multi-scale edge representations better capture structural details and enhance visual fidelity. We further conduct a parameter analysis on the number of inference steps, as shown in Figure 3. As the number of steps increases, robust accuracy shows a slight improvement, while sampling time grows significantly.
Figure 3: Parameter analysis of inference steps.
Conclusion
@@
Acknowledgments
This work was supported by the InnovaLab.ai GPU research incentive for supporting research projects. Research Grant GPU-RTX2025.
Supplementary Material
@@ Appendix A Derivations and Proofs
Proof of Limit Formula
Here, we aim to show that as $t\rightarrow 0$, the difference between the consistency model outputs converges to the adversarial perturbation, i.e., $\boldsymbol{f_{\theta}}(\mathbf{z}_{t}^{a},\varnothing,t)-\boldsymbol{f_{\theta}}(\mathbf{z}_{t},\varnothing,t)\to\boldsymbol{\epsilon}_{a}$.
$\lim_{t\to 0}\boldsymbol{f_{\theta}}(\mathbf{z}_{t}^{a},\varnothing,t)-\boldsymbol{f_{\theta}}(\mathbf{z}_{t},\varnothing,t)$ (19)
@@
$=$ $\lim_{t\to 0}c_{\mathrm{skip}}(t)\sqrt{\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{a}+c_{\mathrm{out}}(t)\left(\frac{\sqrt{\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{a}-\sigma_{t}(\hat{\epsilon}_{\theta}^{a}-\hat{\epsilon}_{\theta})}{\alpha_{t}}\right)$
$=$ $\lim_{t\to 0}c_{\mathrm{skip}}(t)\sqrt{\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{a}$
$=$ $\boldsymbol{\epsilon}_{a}$
Derivation of Equation (10)
In Equation (9), our objective is to select kt such that the adversarial perturbation $\boldsymbol{\epsilon}_{a}$ is effectively removed, given that only the adversarial latent $\mathbf{z}^{a}$ is available at inference. Following (Dhariwal and Nichol 2021), we leverage Bayes’ theorem and the properties of Gaussian distributions to rewrite the sampling formulation as:
@@
$\mathrm{B}=-2\sqrt{\alpha_{t}}\cdot\frac{\tilde{\mathbf{z}}_{t}-(\sqrt{\alpha_{t}}k_{t-1}-k_{t})\boldsymbol{\epsilon}_{a}}{1-\alpha_{t}}$
$\qquad-2\frac{\sqrt{\bar{\alpha}_{t-1}}\mathbf{z}_{0}+(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1})\boldsymbol{\epsilon}_{a}}{1-\bar{\alpha}_{t-1}}$
To achieve our goal, we should remove terms related to $\boldsymbol{\epsilon}_{a}$ in Equation (20), which leads to:
$\frac{\sqrt{\alpha_{t}}(\sqrt{\alpha_{t}}k_{t-1}-k_{t})\boldsymbol{\epsilon}_{a}}{1-\alpha_{t}}=\frac{(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1})\boldsymbol{\epsilon}_{a}}{1-\bar{\alpha}_{t-1}}$ (22)

### v2

@@ 4.3 Adaptive Semantic Enhanced Purification
where the thresholds are calculated using Otsu Otsu (1979) algorithm.
@@
5 Experiments
5.1 Experimental Settings
Table 3: Robust Accuracy ($\%$) results on CelebA under PGD-10. The best results are bolded, and the second-best results are underlined.
@@ 5.1 Experimental Settings
Datasets
We conduct extensive experiments to validate the effectiveness and efficiency of our proposed method across several widely-used datasets, including CIFAR-10 Krizhevsky et al. (2009), ImageNet Deng et al. (2009), and CelebA Liu et al. (2015). CIFAR-10 consists of 60,000 color images of size 32×32 across 10 object classes, representing general-purpose natural scenes. ImageNet is a large-scale visual database with over 14 million human-annotated images spanning more than 20,000 categories. CelebA contains over 200,000 celebrity face images, each annotated with 40 facial attributes and five landmark points.
Table 4: Robust Accuracy ($\%$) on DBLP under Diff-PGD-10 attack $\epsilon=8/255$ on ImageNet.
DiffPure Nie et al. (2022) 53.8 49.4 52.2 42.9 16.6 45.1
OSCP Lei et al. (2025) 59.0 56.5 57.9 49.1 34.1 53.9
DBLP (Ours) 63.0 59.4 60.7 52.4 38.2 58.3
Training Settings
For our pretrained diffusion backbone, we use Stable Diffusion v1.5 Rombach et al. (2022). The distillation process is trained for 20,000 iterations with a batch size of 4, a learning rate of 8e-6, and a 500-step warm-up schedule. For our leapfrog solver $\Psi_{\mathrm{leap}}$, we set $k=20$ in Equation equation 13 and $h=0.8$ in Equation equation 15. During training, adversarial noise is generated using PGD-100 with $\epsilon=4/255$, targeting a ResNet-50 He et al. (2016) classifier.
Evaluation Metrics
We evaluate our approach using multiple metrics: clean accuracy (performance on clean data), robust accuracy (performance under adversarial attack), inference time, and image quality metrics including LPIPS Zhang et al. (2018), PSNR, and SSIM Horé & Ziou (2010).
5.2 Results
CIFAR-10
We first conduct experiments on the CIFAR-10 dataset, evaluating our method under adversarial threats constrained by $\ell_{\infty}$, $\ell_{1}$, and $\ell_{2}$ norms. Since DBLP is trained under $\ell_{\infty}$ attacks, this scenario is considered a seen threat, while the $\ell_{1}$ and $\ell_{2}$ settings are treated as unseen threats. The results are presented in Table 1. Although DBLP belongs to the category of adversarial purification methods, its access to the victim classifier makes it comparable to SOTA adversarial training and purification methods. Adversarial training performs well on seen threats but generalizes poorly to unseen ones, and DiffPure variants offer limited gains. In contrast, DBLP achieves substantially higher robust accuracy on both seen and unseen threats, while preserving strong clean accuracy. It outperforms prior methods by 7.23%, highlighting its robustness, generalization, and efficiency.
ImageNet
Celeb-A
Method runtime (s)
GDMP Wang et al. (2022) $\sim 43$
DiffPure Nie et al. (2022) $\sim 53$
OSCP Lei et al. (2025) $\sim 0.8$
DBLP (Ours) $\sim\textbf{0.2}$
Table 6: Ablation study of adaptive semantic enhancement.
@@ 5.2 Results
w/o Edge Map 74.2 0.1386 0.7409
Edge Map 74.8 0.1172 0.7430
DBLP (Ours) 75.6 0.1012 0.7655
We further validated the effectiveness of our method on a subset of the CelebA-HQ dataset by evaluating it against three representative victim models: ArcFace (AF) Deng et al. (2019), FaceNet (FN) Schroff et al. (2015), and MobileFaceNet (MFN) Chen et al. (2018). Leveraging model weights pretrained on ImageNet, we applied our purification framework to adversarial face images. As shown in Table 3, DBLP significantly enhances purification performance on facial data, demonstrating its robust generalization across image resolutions and domains.
5.3 Image Quality
@@
Beyond correct classification, adversarial purification also seeks to maintain visual fidelity relative to the clean input. As shown in Table 7, DBLP achieves strong image quality across all three metrics, with purified outputs $\mathbf{x}_{\mathrm{pur}}$ closely matching both adversarial $\mathbf{x}_{\mathrm{adv}}$ and clean images $\mathbf{x}$. This highlights DBLP’s superior visual quality. Qualitative results in Figure 2 further confirm its ability to preserve fine-grained details.
Figure 3: Parameter analysis of inference steps.
5.4 Ablation Study
We conduct ablation studies to assess the adaptive semantic enhancement module in DBLP, with results in Table 6. Omitting edge maps leads to a small drop in robust accuracy but a significant decline in image quality. Using pyramid edge maps further improves both metrics, showing that multi-scale edge representations better capture structural details and enhance visual fidelity. We further conduct a parameter analysis on the number of inference steps, as shown in Figure 3. As the number of steps increases, robust accuracy shows a slight improvement, while sampling time grows significantly. For more ablation results, please refer to Appendix A.3.
5.5 Transferability
We further evaluated DBLP under the Diff-PGD attack Xue et al. (2023). The LCM was trained using PGD-generated adversarial noise on ResNet-50 and tested for transfer robustness across diverse architectures, including ResNet-50/152, WideResNet-50-2 Zagoruyko & Komodakis (2016), ConvNeXt-B Liu et al. (2022), ViT-B-16 Kolesnikov et al. (2021), and Swin-B Liu et al. (2021). As shown in Table 4, DBLP consistently outperforms prior SOTA methods under Diff-PGD-10, demonstrating strong cross-architecture robustness.
5.6 Inference Time
@@
6 Conclusion
@@ Appendix A Derivations and Proofs
A.1 Proof of Limit Formula
Here, we aim to show that as $t\rightarrow 0$, the difference between the consistency model outputs converges to the adversarial perturbation, i.e., $\bm{f_{\theta}}(\mathbf{z}_{t}^{a},\varnothing,t)-\bm{f_{\theta}}(\mathbf{z}_{t},\varnothing,t)\to\bm{\epsilon}_{a}$.
$\lim_{t\to 0}\bm{f_{\theta}}(\mathbf{z}_{t}^{a},\varnothing,t)-\bm{f_{\theta}}(\mathbf{z}_{t},\varnothing,t)$ (19)
@@ A.1 Proof of Limit Formula
$=$ $\lim_{t\to 0}c_{\mathrm{skip}}(t)\sqrt{\bar{\alpha}_{t}}\bm{\epsilon}_{a}+c_{\mathrm{out}}(t)\left(\frac{\sqrt{\bar{\alpha}_{t}}\bm{\epsilon}_{a}-\sigma_{t}(\hat{\epsilon}_{\theta}^{a}-\hat{\epsilon}_{\theta})}{\alpha_{t}}\right)$
$=$ $\lim_{t\to 0}c_{\mathrm{skip}}(t)\sqrt{\bar{\alpha}_{t}}\bm{\epsilon}_{a}$
$=$ $\bm{\epsilon}_{a}$
A.2 Derivation of Equation equation 10
In Equation equation 9, our objective is to select kt such that the adversarial perturbation $\bm{\epsilon}_{a}$ is effectively removed, given that only the adversarial latent $\mathbf{z}^{a}$ is available at inference. Following Dhariwal & Nichol (2021), we leverage Bayes’ theorem and the properties of Gaussian distributions to rewrite the sampling formulation as:
@@ A.2 Derivation of Equation equation 10
$\mathrm{B}=-2\sqrt{\alpha_{t}}\cdot\frac{\tilde{\mathbf{z}}_{t}-(\sqrt{\alpha_{t}}k_{t-1}-k_{t})\bm{\epsilon}_{a}}{1-\alpha_{t}}$
$\qquad-2\frac{\sqrt{\bar{\alpha}_{t-1}}\mathbf{z}_{0}+(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1})\bm{\epsilon}_{a}}{1-\bar{\alpha}_{t-1}}$
To achieve our goal, we should remove terms related to $\bm{\epsilon}_{a}$ in Equation equation 20, which leads to:
$\frac{\sqrt{\alpha_{t}}(\sqrt{\alpha_{t}}k_{t-1}-k_{t})\bm{\epsilon}_{a}}{1-\alpha_{t}}=\frac{(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1})\bm{\epsilon}_{a}}{1-\bar{\alpha}_{t-1}}$ (22)
@@
A.3 More Ablation Results
We conducted additional ablation studies to rigorously evaluate the effectiveness of each component in DBLP. Specifically, we ablated Noise Bridge Distillation (NBD) and the Leapfrog ODE solver, comparing them respectively with the consistency distillation loss (CD) of the Latent Consistency Model and the conventional DDIM solver. The distillation loss ablation results, summarized in Table 8, demonstrate that NBD consistently outperforms the traditional CD across all metrics, achieving superior robust accuracy and perceptual image quality. This indicates that, by introducing a noise bridge, NBD more effectively aligns the adversarial noise distribution with the clean data distribution, thereby substantially enhancing both model robustness and the quality of purified images.
Table 8: Ablation study on different distillation loss.
Distillation Loss Robust Accuracy LPIPS SSIM
w/o 65.1 0.1857 0.7260
CD 73.5 0.1337 0.7492
NBD 75.6 0.1012 0.7655
In the ODE solver ablation, the Leapfrog solver exhibits remarkable performance and efficiency, surpassing the DDIM solver. These results confirm that the Leapfrog solver’s distinctive update mechanism enables higher computational efficiency without compromising purification quality.
Table 9: Ablation study on different ODE solvers.
ODE Solver Robust Accuracy LPIPS Time
DDIM 75.40 0.1029 0.2670
Leapfrog 75.60 0.1012 0.2315